# The Binary Fraction of Stars in the Dwarf Galaxy Ursa Minor via Dark Energy Spectroscopic Instrument

### v1

@@ 6 Conclusion
We also divide member stars in UMi according to the projected radius to the center, and after resampling them to have the same metallicity distributions, we find that stars in the denser central region have lower binary fractions ($0.33^{+0.30}_{-0.20}$) than the outer part ($1.00^{+0.00}_{-0.32}$) with the 2017ApJS..230...15M model.
@@
Future surveys with more exposures crossing longer time baselines and higher signal-to-noise spectroscopy will be crucial in further verifying these findings. Extended observations of the binary population in UMi, with a better sampling of orbital periods and a more precise understanding of the binary fraction, will allow for a more thorough exploration of the mechanisms that drive binary survival in metal-poor environments. Moreover, similar studies in other dwarf galaxies within the Milky Way will help to determine whether the metallicity trends observed in UMi are applicable more broadly across different environments. This will ultimately deepen our understanding of binary star evolution, especially in the context of early galaxy formation and the complex interplay among chemical evolution, star formation, and dynamical interaction.
@@
The results presented in this paper can be accessed at https://zenodo.org/records/16827506, which contains all data points for the figures.
@@
This work is supported by NSFC (12273021, 12022307, 12573022, 12595312) and the National Key R&D Program of China (2023YFA1605600, 2023YFA1605601). We thank the sponsorship from Yangyang Development Fund. The computations of this work are carried on the Gravity supercomputer at the Department of Astronomy, Shanghai Jiao Tong University and on the National Energy Research Scientific Computing Center (NERSC). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. We thank all contributors to this work for their valuable input and support. We thank the DESI operation team, Stephen Bailey, Anand Raichoor, and Edward Schlafly for their contributions to the maintenance of the tertiary programs in the DESI MWS. We thank our DESI publication handlers, Ting Tan and Alejandro Avilés, for their assistance. We are grateful to the organizers and participants of the International Astronomical Union (IAU) Symposium 398 for the insightful discussions and constructive feedback that helped shape and improve this study. For the purpose of open access, the author has applied a Creative Commons Attribution ($C$ BY) licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ VI Conclusion
We also divide member stars in UMi according to the projected radius to the center, and after resampling them to have the same metallicity distributions, we find that stars in the denser central region have lower binary fractions ($0.33^{+0.30}_{-0.20}$) than the outer part ($1.00^{+0.00}_{-0.32}$) with the MS17 model.
@@
Future surveys with more exposures crossing longer time baselines and higher signal-to-noise spectroscopy will be crucial in further verifying these findings. Although the present analysis achieves high coverage of confirmed member candidates, matching over 90% of the current target list, only about 50% of these stars are included in the final analysis due to insufficient epoch coverage or large LOSV uncertainties. Extended observations that provide additional epochs and improved velocity precision will therefore enhance sensitivity to longer-period binaries and enable more robust constraints on the binary population in UMi. With improved sampling of orbital periods and tighter constraints on the binary fraction, future studies will be better positioned to explore the physical mechanisms governing binary survival in metal-poor environments. Moreover, similar studies in other dwarf galaxies within the Milky Way will help to determine whether the metallicity trends observed in UMi are applicable more broadly across different environments. This will ultimately deepen our understanding of binary star evolution, especially in the context of early galaxy formation and the complex interplay among chemical evolution, star formation, and dynamical interaction.
@@
The results presented in this paper can be accessed at https://doi.org/10.5281/zenodo.16827505, which contains all data points for the figures.
@@
This work is supported by NSFC (12573022, 12595312, 12273021), the National Key R&D Program of China (2023YFA1605600, 2023YFA1605601), 111 project (No. B20019), and the Office of Science and Technology, Shanghai Municipal Government (grant Nos. 24DX1400100, ZJ2023-ZD-001). We thank the sponsorship from Yangyang Development Fund. The computations of this work are carried on the National Energy Research Scientific Computing Center (NERSC) and the Gravity supercomputer at the Department of Astronomy, Shanghai Jiao Tong University. SK acknowledges support from the Science & Technology Facilities Council (STFC) grant ST/Y001001/1. This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. We thank all contributors to this work for their valuable input and support. We thank the DESI operation team, Stephen Bailey, Anand Raichoor, and Edward Schlafly for their contributions to the maintenance of the tertiary programs in the DESI MWS. We thank our DESI publication handlers, Ting Tan and Alejandro Avilés, for their assistance. We are grateful to the organizers and participants of the International Astronomical Union (IAU) Symposium 398 for the insightful discussions and constructive feedback that helped shape and improve this study. For the purpose of open access, the author has applied a Creative Commons Attribution ($C$ BY) licence to any Author Accepted Manuscript version arising from this submission.